In a survey, the planning value for the population proportions P* = 0.5. How large a sample should be taken to provide a 90% confidence interval with a margin of error of 0.07?
Please keep two decimal places of your answer.

Answer by Boreal(15235)

margin of error is z*sqrt(p*(1-p)/n)=0.07 and z(0.95)=1.645 for a CI of 90%
square both sides and 1.645^2*0.25/n=0.0049
2.706025*0.25=0.0049n
n=138.06